OpenStudy (anonymous):

What is the LCD for the fractions 3/16 and 11/24? A. 24 B. 40 C. 48 D. 384

OpenStudy (anonymous):

LCD is the Lowest Common Multiple in the Denominator. so, i'm looking at two numbers 16, and 24 i'm going to break them down into their smallest prime numbers 16 = 4*4 or = 2*2*2*2 24 = 12*2 or = 6*2*2 or = 2*3*2*2 now i count how many times 2 shows up in each number. in 16 it shows up 4 times in 24 it show ups 3 times. now i count up how many times 3 shows up in each number in 16 it shows up 0 times in 24 it shows up 1 time. now I take the max number 2 occurs and 3 occurs, that is four 2's and one 3 to get my LCD: 2*2*2*2*3

OpenStudy (anonymous):

48